OpenStudy (anonymous):

Find the limit of (1-x)^(lnx) as x--> 1

OpenStudy (lgbasallote):

let y = (1-x)^lnx lny = lnx ln(1-x)

OpenStudy (anonymous):

yeah that is right

OpenStudy (anonymous):

then rewrite as \[\frac{\ln(1-x)}{\frac{1}{\ln(x)}}\] and use l'hopital

OpenStudy (lgbasallote):

ahhh LH...a very frustrating method involving so much manipulation and circling

OpenStudy (anonymous):

LH is frustrating?! it's makes your life easy.

OpenStudy (anonymous):

thanks guys you nailed it =]

OpenStudy (lgbasallote):

it's frustating when it circles

OpenStudy (anonymous):

if it circles then you are missing some thing

OpenStudy (anonymous):

derivative of numerator is \[-\frac{1}{1-x}\] derivative of denominator is \[-\frac{1}{x\ln^2(x)}\] ratio is \[\frac{x\ln^2(x)}{1-x}\]

OpenStudy (lgbasallote):

also... 0/0 is easy.... i hate (inf - inf) 1^inf and something

OpenStudy (anonymous):

x goes to one, get 0/1 so not so bad. limit of the log is 0, limit of the original one is \[e^0=1\] unless i made a mistake

OpenStudy (anonymous):

yes you are right 1 is the answer ! =]